# Majority-Logic Decoding of Binary Locally Recoverable Codes: A Probabilistic Analysis

### v1

@@ I Introduction
Locally repairable codes (LRCs) are a class of error-correcting codes designed to enable the recovery of individual code symbols by accessing only a small subset of other symbols, known as local recovery sets (also referred to as repair groups). This property, formalized through the notions of locality and availability, makes LRCs particularly attractive for distributed storage systems, where node failures lead to data loss modeled as erasures [10]. By providing multiple, disjoint recovery sets for each symbol, LRCs allow local repair with low bandwidth and latency overheads, and have therefore received significant attention in recent years.
@@
Although LRCs were originally introduced to combat erasures, the existence of multiple recovery sets naturally suggests their applicability to error correction as well. In practical storage and memory systems, failures are often better modeled as random errors—such as bit flips—rather than erasures, reflecting data corruption at the level of memory elements [25]. Such bit flips arise from a variety of unavoidable physical effects and remain a major concern even in highly reliable systems, including safety-critical applications [2]. When errors are present, corrupted symbols may invalidate some recovery sets, thereby affecting the reliability of local decoding. This observation motivates the use of majority-logic decoding (MLD), a classical decoding paradigm first introduced for Reed–Muller codes [28, 20]. Under MLD, each symbol is recovered by taking a majority vote over multiple recovery sets, where each set provides an independent linear equation for the target symbol. Decoding succeeds as long as a majority (i.e., at least half) of these equations remain uncorrupted [18].
@@
Majority-logic decoding is a simple, hard-decision decoding method with low complexity and extremely small decoding delay, which historically made it attractive for applications such as deep-space communication and memory protection [15, 16, 9]. Despite major advances in decoding algorithms and hardware capabilities, there has been a renewed interest in such low-complexity decoders for latency-critical applications, where hard-decision algorithms are often preferred over soft-decision alternatives [4]. Recently, MLD has also been connected to combinatorial designs, and to the concept of Service Rate Region, a metric that quantifies the efficiency of linear codes for optimizing data access [17]. Moreover, MLD is not a bounded-distance decoder and is known to frequently correct more errors than its guaranteed radius [6]. For certain code families, such as Grassmann codes, MLD remains the only known method enabling efficient and practical decoding [3]. Codes that admit such decoding algorithms are referred to as majority-logic decodable codes [23].
@@ II-A Locally Recoverable Codes with Availability
In this write-up, we focus on a class of binary linear LRCs $\mathscr{C}\subseteq\mathbb{F}_{2}^{n}$ in which: (1) all recovery sets have the same size $r$, where $r$ is fixed, and (2) the number of disjoint recovery sets $t$ grows with the code length $n$ (as in the case of Simplex codes). We justify that the assumption of constant recovery set size (1) is both valid and reasonable, as used commonly in the literature, for example, in [29]. Indeed, this assumption not only helps simplify our analysis but is practically solid. If a recovery set of size $s<r$ exists, it can always be formally treated as a set of size $r$ by add into it other rsr-s available symbols among the remaining. Since these extra objects do not participate in the actual recovery process, the (pairwise) disjointness of the recovery sets is preserved. Moreover, in hardware/VLSI (Very-Large-Scale Integration) systems, it is significantly more efficient to design, implement, and maintain circuits with a uniform number of inputs rather than supporting varying input sizes [11], especially as the number of required circuits scales. Such codes are often referred to as LRCs with large availability and have attracted significant practical interest [13]. Explicit constructions of binary linear LRCs with fixed locality and availability growing with $n$ have been presented, for example, in [13] (linear growth) and [22, 29]. Authors in [27] presented a construction of a class of LRC with scaling locality and availability that required large field sizes. Authors in [30] presented a construction of a class of binary code that achieves arbitrary locality and availability. Interested readers are referred to [26] for a comprehensive and detailed review.
@@ II-B Majority-Logic Decoding
Thus, $\{c_{2},c_{3}\}$, $\{c_{4},c_{5}\}$, and $\{c_{6},c_{7}\}$ form three independent recovery sets for $c_{1}$.
Suppose the received vector $\mathbf{u}$ differs from the transmitted codeword $\mathbf{c}$ due to channel errors. Since the recovery sets in (1) are pairwise disjoint, each bit error affects at most one recovery set. For example, if only $u_{2}$ is flipped while all other symbols are received correctly, then the estimate $u_{2}+u_{3}$ disagrees with $u_{4}+u_{5}$ and $u_{6}+u_{7}$. A majority vote among
@@
In contrast, an error pattern of the same weight affecting two distinct recovery sets (e.g., errors at $u_{2}$ and $u_{4}$) leads to an incorrect majority vote. This illustrates that a recovery set produces an incorrect vote if and only if it contains an odd number of errors.
In general, a local recovery set containing an even number of errors yields a correct vote, whereas a recovery set containing an odd number of errors yields an incorrect vote. Thus, the correctness of each vote depends only on the parity of the number of errors within the recovery set, not on the total number of errors. Consequently, majority-logic decoding can successfully correct certain error patterns that conventional bounded-distance decoding methods cannot, despite relying solely on local parity constraints. This observation motivates our probabilistic analysis of majority-logic decoding for $(r,t)_{a}$ LRCs over channels with independent bit flips and erasures, in terms of BER, BLER, and the number of correctable errors and erasures.
@@ III-A Deterministic versus Probabilistic Perspectives
However, such worst-case bounds are often pessimistic and not particularly informative in practice. In contrast, a more common and realistic scenario is one in which errors and erasures occur randomly [21]. Under such stochastic behavior, errors often affect multiple symbols within the same recovery set, leaving other recovery sets completely intact. This sharply contrasts with the adversarial scenario, where errors are distributed to corrupt as many recovery sets as possible. Consequently, the number of corrupted recovery sets is often significantly smaller than the total number of raw errors, allowing successful decoding even beyond worst-case guarantees. This explains why MLD frequently corrects more than $\lfloor(t-1)/2\rfloor$ errors in practice and motivates the probabilistic analysis that follows.
@@ III-B Binary Symmetric Channel ($\mathrm{BSC}$)
$P^{\mathrm{BSC}}_{\mathrm{fail,bit}}$ $\leq\bigl(2\sqrt{a(1-a)}\bigr)^{t}=\big(1-(1-2p_{f})^{2r}\big)^{t/2},$
$\quad\text{where}\quad a=\frac{1+(1-2p_{f})^{r}}{2}.$
@@
For a given recovery set $R^{x}$ of size $r$, let
@@
Then $\Pr[\xi_{x}=1]=C$ and $\Pr[\xi_{x}=-1]=W$, consequently
$\mathbb{E}[\xi_{x}]=C-W=(1-2p_{f})^{r}=y^{r},\qquad\mathrm{Var}(\xi_{x})=1-y^{2r},$
where $y\triangleq 1-2p_{f}$. MLD fails when the total number of wrong votes exceeds the correct ones, i.e.,
@@
The result follows from the standard exponential moment method (see, e.g., [5, Ch. 2]). For any $\lambda<0$, Markov’s inequality applied to the random variable $e^{\lambda S_{t}}$ implies
@@
Consequently, $\Pr(S_{t}\leq 0)\approx\mathrm{e}^{-t\cdot D_{\mathrm{KL}}(1/2\,\|\,a)}=\left(2\sqrt{a(1-a)}\right)^{t}$. This confirms that both the base and the linear dependence on $t$ are asymptotically optimal. Note that although the base of the exponent is optimal, the bound can be further refined by incorporating the polynomial pre-factor established by the Bahadur-Rao theorem [1].
Remark 2.
Theorem 1 does not assume how $t$ scales with $n$. In particular, the bound holds uniformly for all values of $n$ and $t$, regardless of whether $t$ is fixed, grows slowly, or grows unbounded with $n$.
On the other hand, it follows from (2) that the quantity CC—the probability that a given recovery set produces a correct vote—is a decreasing function of the locality $r$ for a fixed bit-flip probability $p_{f}\in[0,\,0.5)$. This behavior may appear counterintuitive, since $C$ depends only on the parity (even or odd) of the number of bit flips within the recovery set. This implies that, when the availability $t$ is fixed, smaller recovery sets yield a higher probability of a correct vote.
@@
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all error patterns of weight $w$ are correctable in the asymptotic regime of $n$.
@@
$p\triangleq\frac{1}{2}\!\left(1-\sqrt[2r]{\frac{c\log n}{t}}\right),\qquad x\triangleq(1-2p)^{2r}=\frac{c\log n}{t},$
@@
$=\exp\!\left(\left(1-\frac{c}{2}\right)\log n-\frac{t}{2}\varepsilon\right)\ (\text{since }tx=c\log n.)$
@@
By the classical lower bound on binomial probabilities [19, Ch. 10, Lem. 7],
@@
$=\Theta(n^{-1/2})=\exp\!\left(-\frac{1}{2}\log n+o(1)\right).$
@@
Substituting the upper bound for $P_{\mathrm{fail,block}}^{\mathrm{BSC}}$ and the lower bound for $P(A_{m})$,
$\frac{|B_{m}|}{|A_{m}|}\leq\exp\!\left(\frac{3}{2}\log n-\frac{c}{2}\log n-\frac{t}{2}\varepsilon-o(1)\right).$
@@
Remark 3.
The constant $c$ in  (4) and (III-C) captures a trade-off between the weight of the correctable error patterns and the rate at which the fraction of uncorrectable patterns decays. Choosing a smaller value of $c$ increases the threshold weight $w$ in (4), allowing MLD to correct error patterns of higher weight. However, this comes at the cost of a slower decay in the upper bound (III-C) on the fraction of uncorrectable patterns. Conversely, larger values of $c$ yield faster decay of this fraction but guarantee correctability only for error patterns of smaller weight.
@@
Remark 5 (Numerical example).
@@ III-C Binary Erasure Channel ($\mathrm{BEC}$)
Since $0\leq 1-(1-p_{e})^{r}<1$, the probability of decoding failure decays as $t$ grows. This observation leads to the following interesting result.
@@
Let $c\geq\frac{3}{2}$ be any constant. For a binary $(r,t)_{a}$-LRC with availability $t=t(n)$ growing with $n$, all but a vanishing fraction of erasure patterns of weight
@@
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all erasure patterns of weight $w$ are correctable in the asymptotic regime of $n$.
@@
Moreover, by the classical binomial lower bound [19, Ch. 10, Lem. 7],

### v2

@@ I Introduction
Locally repairable codes (LRCs) are a class of error-correcting codes designed to enable the recovery of individual code symbols by accessing only a small subset of other symbols, known as local recovery sets (also referred to as repair groups). This property, formalized through the notions of locality and availability, makes LRCs particularly attractive for distributed storage systems, where node failures lead to data loss modeled as erasures [11]. By providing multiple, disjoint recovery sets for each symbol, LRCs allow local repair with low bandwidth and latency overheads, and have therefore received significant attention in recent years, with applications extending beyond distributed storage [15].
@@
Although LRCs were originally introduced to combat erasures, the presence of multiple recovery sets also suggests potential robustness to errors. In practical storage and memory systems, failures are often better modeled as random errors—such as bit flips—rather than erasures, reflecting data corruption at the level of memory elements [31]. Such bit flips may arise from unavoidable physical effects and remain a concern even in highly reliable systems, including safety-critical applications [1]. When errors are present, corrupted symbols may invalidate some recovery sets, thereby degrading the reliability of local recovery. This observation motivates the use of majority-logic decoding (MLD), a classical decoding paradigm first introduced for Reed–Muller (RM) codes [34, 26]. Under MLD, each recovery set produces a local estimate (a vote) for the target symbol, and the decoder outputs the majority of these votes (with an arbitrary tie-breaking rule when needed) [24]. In the binary case, each local estimate is a parity computation and can be implemented using XOR operations. More broadly, highly efficient erasure-coding families such as Tornado codes and fountain codes (e.g., LT and Raptor codes) rely on sparse parity constraints and iterative decoding procedures whose computations are mostly just XOR operations, enabling low-complexity (often linear- or near-linear-time) encoding and decoding on erasure channels [21, 20, 36].
@@
Majority-logic decoding is a simple, hard-decision decoding method with low complexity and extremely small decoding delay, which historically made it attractive for applications such as deep-space communication and memory protection [18, 19, 10]. Despite major advances in decoding algorithms and hardware capabilities, there has been a renewed interest in such low-complexity decoders for latency-critical applications, where hard-decision algorithms are often preferred over soft-decision alternatives [3]. Recently, MLD has also been connected to combinatorial designs, and to the concept of Service Rate Region, a metric that quantifies the efficiency of linear codes for optimizing data access [23]. Moreover, MLD is not a bounded-distance decoder and is known to frequently correct more errors than other bounded-distance decoders [6]. For instance, over long binary RM codes of fixed rate, MLD can correct most error patterns of weight up to $d\log d/4$, where $d$ denotes the minimum distance [17]. For certain code families, such as Grassmann codes, MLD remains the only known method enabling efficient and practical decoding [2]. Codes that admit such decoding algorithms are referred to as majority-logic decodable codes [29].
@@
We note that LRCs constitute a broadly defined class of codes characterized primarily by their local recovery properties—namely, locality and, in some cases, availability—rather than by a specific algebraic construction such as a generator or parity-check matrix. This distinguishes LRCs from classical code families, for which decoding algorithms are often tailored to explicit algebraic structures (e.g., the Berlekamp–Welch algorithm for Reed–Solomon codes or the Berlekamp–Massey decoder for BCH codes). As a result, there is currently no canonical, structure-specific decoding algorithm for general LRCs. In this sense, MLD emerges as a natural and broadly applicable decoding framework for LRCs, as it relies solely on the presence of multiple local recovery sets inherent to the codes’ definition.
@@ II-A Locally Recoverable Codes with Availability
In this write-up, we focus on a class of binary linear LRCs $\mathscr{C}\subseteq\mathbb{F}_{2}^{n}$ in which: (1) all recovery sets have the same size $r$, where $r$ is fixed, and (2) the number of disjoint recovery sets $t$ grows with the code length $n$ (as in the case of Simplex codes). We justify that the assumption of constant recovery set size (1) is both valid and reasonable, as used commonly in the literature, for example, in [37, 16]. Indeed, this assumption not only helps simplify our analysis but is practically solid. If a recovery set of size $s<r$ exists, it can always be formally treated as a set of size $r$ by add into it other rsr-s available symbols among the remaining. Since these extra objects do not participate in the actual recovery process, the (pairwise) disjointness of the recovery sets is preserved. Moreover, in hardware/VLSI (Very-Large-Scale Integration) systems, it is significantly more efficient to design, implement, and maintain circuits with a uniform number of inputs rather than supporting varying input sizes [12], especially as the number of required circuits scales. Such codes are often referred to as LRCs with large availability and have attracted significant practical interest [14]. Explicit constructions of binary linear LRCs with fixed locality and availability growing with $n$ have been presented, for example, in [14] (linear growth) and [28, 37]. Authors in [33] presented a construction of a class of LRC with scaling locality and availability that required large field sizes. Authors in [38] presented a construction of a class of binary code that achieves arbitrary locality and availability. Interested readers are referred to [32] for a comprehensive and detailed review.
@@ II-B Majority-Logic Decoding
Equivalently, $R^{1,1}=\{2,\,3\},\,R^{2,1}=\{4,\,5\},\,R^{3,1}=\{6,\,7\}$. Suppose the received vector $\mathbf{u}$ differs from the transmitted codeword $\mathbf{c}$ due to channel errors. Since the recovery sets in (1) are pairwise disjoint, each bit error affects at most one recovery set. For example, if only $u_{2}$ is flipped while all other symbols are received correctly, then the estimate $u_{2}+u_{3}$ disagrees with $u_{4}+u_{5}$ and $u_{6}+u_{7}$. A majority vote among
@@
In contrast, an error pattern of the same weight affecting two distinct recovery sets (e.g., errors at $u_{2}$ and $u_{4}$, i.e., $\mathbf{e}=0101000$) causes the majority vote to output an incorrect estimate of $c_{1}$. This highlights that a given recovery set produces an incorrect vote if and only if it contains an odd number of errors.
Such an error pattern contributes to the bit error rate (BER) because the decoded value of $c_{1}$ is incorrect, and it also contributes to the block error rate (BLER) under symbol-wise decoding, since a block is declared correct only when all symbols are decoded correctly.
In general, a local recovery set containing an even number of errors yields a correct vote, whereas a recovery set containing an odd number of errors yields an incorrect vote. Thus, the correctness of each vote depends only on the parity of the number of errors within the recovery set, not on the total number of errors. Consequently, majority-logic decoding can successfully correct certain error patterns that conventional bounded-distance decoding methods cannot, despite relying solely on local parity constraints. This observation motivates our probabilistic analysis of majority-logic decoding for $(r,t)_{a}$ LRCs over channels with independent bit flips (also called crossovers) and erasures, in terms of BER, BLER, and the number of correctable errors and erasures. On a related note, another example of codes where error correction capability depends on error pattern and not on the number of errors itself is Single parity-check Product codes [5].
Following [9], to define their error-correcting performance, we use the following definition. Given an infinite sequence of codes $A_{i}(n_{i},d_{i})$, we say that a decoding algorithm $\Psi$ has a threshold sequence $\delta_{i}$ and a residual sequence $\epsilon_{i}$ if for $n_{i}\rightarrow\infty$, $\Psi$ correctly decodes all but a vanishing fraction of error patterns of weight $\delta_{i}(1-\epsilon_{i})$ or less.
@@ III-A Deterministic versus Probabilistic Perspectives
However, such worst-case bounds are often pessimistic and not particularly informative in practice. In many systems, bursty or dependent error patterns can be mitigated by interleaving, which spreads error bursts across code symbols and can make the resulting error process appear approximately independent at the decoder [35]. In contrast, a more common and realistic scenario is one in which errors and erasures occur randomly [27]. Under such stochastic behavior, errors often affect multiple symbols within the same recovery set, leaving other recovery sets completely intact. This sharply contrasts with the adversarial scenario, where errors are distributed to corrupt as many recovery sets as possible. Consequently, the number of corrupted recovery sets is often significantly smaller than the total number of raw errors, allowing successful decoding even beyond worst-case guarantees. This explains why MLD frequently corrects more than $\lfloor(t-1)/2\rfloor$ errors in practice and motivates the probabilistic analysis that follows.
@@ III-B Binary Symmetric Channel ($\mathrm{BSC}$)
$P^{\mathrm{BSC}}_{\mathrm{fail,bit}}$ $\leq\bigl(2\sqrt{a(1-a)}\bigr)^{t}=\big(1-(1-2p_{f})^{2r}\big)^{t/2},\quad\text{where}\quad a=\frac{1+(1-2p_{f})^{r}}{2}.$
@@
For a symbol $x$ and any of its associated recovery sets $R^{x}$ of size $r$, let
@@
Then $\Pr[\xi_{x}=1]=C$ and $\Pr[\xi_{x}=-1]=W$. MLD fails when the total number of wrong votes exceeds the correct ones, i.e.,
@@
The result follows from the standard exponential moment method (see, e.g., [4, Ch. 2]). For any $\lambda<0$, Markov’s inequality applied to the random variable $e^{\lambda S_{t}}$ implies
@@
Consequently, $\Pr(S_{t}\leq 0)\approx\mathrm{e}^{-t\cdot D_{\mathrm{KL}}(1/2\,\|\,a)}=\left(2\sqrt{a(1-a)}\right)^{t}$. This confirms that both the base and the linear dependence on $t$ are asymptotically optimal.
Theorem 1 makes no assumption on how the availability $t$ scales with the blocklength $n$. In particular, the bound holds uniformly for all $n$ and $t$, whether $t$ is fixed, grows slowly, or grows unbounded with $n$, and it implies an exponential decay (in $t$) of the upper bound on the bit decoding failure probability. On the other hand, it follows from (2) that the probability $C$ that a given recovery set produces a correct vote is a decreasing function of the locality $r$ for any fixed bit-flip probability $p_{f}\in[0,0.5)$; equivalently, smaller recovery sets are more likely to yield a correct vote.
Taken together, these observations reveal a basic contrast. From the perspective of a target symbol, having more recovery sets (larger $t$) provides more independent votes and improves reliability. From the perspective of each recovery equation (parity constraint), smaller locality (smaller $r$) improves the reliability of the individual vote. Interpreting the recovery structure through a bipartite Tanner-like graph, these preferences resemble the tension between variable-node connectivity and check-node degree observed in iterative decoding of LDPC codes (e.g., under belief propagation) [22].
@@
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all error patterns of weight $w$ are correctable in the asymptotic regime of $n$. In other words, long RM codes of growing availability can be decoded with complexity $O(nt(n))$ and decoding threshold $\delta=\dfrac{n}{2},\,\epsilon=\sqrt[2r]{\frac{c\log n}{t(n)}}$.
@@
$p\triangleq\frac{1}{2}\!\left(1-\sqrt[2r]{\frac{c\log n}{t}}\right),\qquad x\triangleq(1-2p)^{2r}=\frac{c\log n}{t},$ (4)
@@
$=\exp\!\left(\left(1-\frac{c}{2}\right)\log n-\frac{t}{2}\varepsilon\right)\ (\text{since }tx=c\log n.)$ (5)
@@
By the classical lower bound on binomial probabilities [25, Ch. 10, Lem. 7],
@@
$=\Theta(n^{-1/2})=\exp\!\left(-\frac{1}{2}\log n+o(1)\right).$ (6)
@@
Substituting the upper bound for $P_{\mathrm{fail,block}}^{\mathrm{BSC}}$ in (III-B) and the lower bound for $P(A_{m})$ in (III-B), one has
$\frac{|B_{m}|}{|A_{m}|}\leq\exp\!\left(\frac{3}{2}\log n-\frac{c}{2}\log n-\frac{t}{2}\varepsilon-o(1)\right).$ (7)
@@
[…]
Theorem 3 can be recast as the following.
Corollary 1.
Long RM codes can be decoded with vanishing output (block) error probability and complexity $O(nt(n))$ on a BSC with bit flip (or crossover) probability $(1-\epsilon)/2$.
One notes that the constant $c$ in (3) and (7) captures a trade-off between the weight of the correctable error patterns and the rate at which the fraction of uncorrectable patterns decays. Choosing a smaller value of $c$ increases the threshold weight $w$ in (3), allowing MLD to correct error patterns of higher weight. However, this comes at the cost of a slower decay in the upper bound (7) on the fraction of uncorrectable patterns. Conversely, larger values of $c$ yield faster decay of this fraction but guarantee correctability only for error patterns of smaller weight.
@@
Example 2 (Numerical example).
@@ III-C Binary Erasure Channel ($\mathrm{BEC}$)
Since $0\leq 1-(1-p_{e})^{r}<1$, the probability of decoding failure decays as $t$ grows. This observation leads to the following result.
@@
Let $c\geq\dfrac{3}{2}$ be any constant. For a binary $(r,t)_{a}$-LRC with availability $t=t(n)$ growing with $n$, all but a vanishing fraction of erasure patterns of weight
@@
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all erasure patterns of weight $w$ are correctable in the asymptotic regime of $n$. In other words, long RM codes of growing availability can be decoded with complexity $O(nt(n))$ and decoding threshold $\delta=n,\,\epsilon=\sqrt[r]{\frac{c\log n}{t(n)}}$.
@@
Moreover, by the classical binomial lower bound [25, Ch. 10, Lem. 7],